OpenStudy (anonymous):

If f(x) = x + 2, find f(3).

OpenStudy (anonymous):

\[f(x)=x+2\] \[f(3)=...+2\]

OpenStudy (unklerhaukus):

\[f(x) = (x) + 2\]\[f(3) = (3) + 2\]

OpenStudy (anonymous):

does it make sense?

OpenStudy (anonymous):

not really ?

OpenStudy (anonymous):

\[x\] can be any number and in this specific problem we want \[x\] to equal 3

OpenStudy (anonymous):

what if x was 1 @delgado1996 what would you do then?

OpenStudy (anonymous):

is it 1

OpenStudy (anonymous):

\[f(x)= x+2\] \[f(3)= 3+2=5\] \[f(2)= 2+2=4\] \[f(1)= 1+2=3\]

OpenStudy (unklerhaukus):